# Efficient 3D Instance Mapping and Localization with Neural Fields https://gtangg12.github.io/iML

### v1

@@ Abstract
We tackle the problem of learning an implicit scene representation for 3D instance segmentation from a sequence of posed RGB images. Towards this, we introduce 3DIML, a novel framework that efficiently learns a label field that may be rendered from novel viewpoints to produce view-consistent instance segmentation masks. 3DIML significantly improves upon training and inference runtimes of existing implicit scene representation based methods. Opposed to prior art that optimizes a neural field in a self-supervised manner, requiring complicated training procedures and loss function design, 3DIML leverages a two-phase process. The first phase, InstanceMap, takes as input 2D segmentation masks of the image sequence generated by a frontend instance segmentation model, and associates corresponding masks across images to 3D labels. These almost view-consistent pseudolabel masks are then used in the second phase, InstanceLift, to supervise the training of a neural label field, which interpolates regions missed by InstanceMap and resolves ambiguities. Additionally, we introduce InstanceLoc, which enables near realtime localization of instance masks given a trained label field and an off-the-shelf image segmentation model by fusing outputs from both. We evaluate 3DIML on sequences from the Replica and ScanNet datasets and demonstrate 3DIML’s effectiveness under mild assumptions for the image sequences. We achieve a 14-24$\times$ speedup over existing implicit scene representation methods with comparable quality, showcasing its potential to facilitate faster and more effective 3D scene understanding.
@@ I Introduction
  • An overall training time improvement of 14-24$\times$ over prior art, over scenes comprising hundreds of objects, on a single GPU (NVIDIA RTX 3090)
@@ IV-D Results
Comparison with Panoptic and Contrastive Lifting: Table II shows the Scene Level Panoptic Quality for 3DIML and other methods on Replica-vMap sequences subsampled by 10 (200 frames). We observe 3DIML approaches Panoptic and Contrastive Lifting in performance despite not relying on updating or querying explicit/implicit world models for much of the pipeline, which makes it much faster. Figure 4 shows a comparison between the instances identified by Panoptic Lifting compared to 3DIML.
We benchmark all runtimes using a single RTX 3090 post-mask generation. Specifically, Panoptic Lifting requires on average 5.7 hours of training over all scans, with min and max of 3.6 and 6.6 hours, respectively, since its runtime is dependent on the number of objects. Contrastive Lifting takes around 3.5 hours on average while 3DIML runs under 20 minutes (14.5 minutes on average) for all scans. Note several components of 3DIML can be easily parallelized, such as dense descriptors extraction using LoFTR and label merging. The runtime of our method is dependent on the number of correspondences produced by LofTR, which doesn’t change for different frontend segmentation models, and we observe similar runtimes for other experiments.
@@
Method Time Panoptic lifting [5] 5.7 hours Contrastive lift [6] 3.6 hours InstanceMap (w/ LoFTR) 6.2 minutes InstanceMap 2.1 minutes InstanceLift (train) 5.6 minutes InstanceLift (refine) 33.2 seconds 3DIML (total) 14.5 minutes
@@
Grounded SAM Table III shows our results for lifting GroundedSAM masks for Replica-vMap. From Figure 5 we see that InstanceLift is effective at interpolating labels missed by InstanceMap and resolving ambiguities produced by GroundedSAM  [^1] GroundedSAM produces lower quality frontend masks than SAM due to prompting using bounding boxes instead of a point grid.. Figure 7 shows that InstanceMap and 3DIML are robust to large viewpoint changes and as well as duplicate objects, assuming nice scans, that is enough context for NetVLAD and LoFTR to somewhat distinguish between them. This is achieved by efficiently relying on implicit scene representation methods only at critical junctions, which gives 3DIML a large edge in runtime compared to Panoptic and Contrastive Lifting. Table IV and Fig. 6 illustrate our performance on ScanNet [21].
@@
Figure 7: Qualitative results on office3 and room1 from the Replica Vmap split [20]. Both InstanceMap and InstanceLift are able to maintain quality and consistency over the image sequence despite duplicate objects due to sufficient image context overlap across the sequence.

### v2

@@ Abstract
We tackle the problem of learning an implicit scene representation for 3D instance segmentation from a sequence of posed RGB images. Towards this, we introduce 3DIML, a novel framework that efficiently learns a label field that may be rendered from novel viewpoints to produce view-consistent instance segmentation masks. 3DIML significantly improves upon training and inference runtimes of existing implicit scene representation based methods. Opposed to prior art that optimizes a neural field in a self-supervised manner, requiring complicated training procedures and loss function design, 3DIML leverages a two-phase process. The first phase, InstanceMap, takes as input 2D segmentation masks of the image sequence generated by a frontend instance segmentation model, and associates corresponding masks across images to 3D labels. These almost view-consistent pseudolabel masks are then used in the second phase, InstanceLift, to supervise the training of a neural label field, which interpolates regions missed by InstanceMap and resolves ambiguities. Additionally, we introduce InstanceLoc, which enables near realtime localization of instance masks given a trained label field and an off-the-shelf image segmentation model by fusing outputs from both. We evaluate 3DIML on sequences from the Replica and ScanNet datasets and demonstrate 3DIML’s effectiveness under mild assumptions for the image sequences. We achieve a large practical speedup over existing implicit scene representation methods with comparable quality, showcasing its potential to facilitate faster and more effective 3D scene understanding.
@@ I Introduction
  • An overall practical runtime improvement of 14-24$\times$ over prior art benchmarked on a single GPU (NVIDIA RTX 3090)
@@ IV-D Results
Comparison with Panoptic and Contrastive Lifting: Table II shows the Scene Level Panoptic Quality for 3DIML and other methods on Replica-vMap sequences subsampled by 10 (200 frames). We observe 3DIML approaches Panoptic Lifting in performance while achieving a much larger practical runtime (considering implementation) than Panoptic and Contrastive Lifting. Intuitively, this is achieved by efficiently relying on implicit scene representation methods only at critical junctions i.e. post InstanceMap, greatly reducing the number of training iterations of the neural field (25x less). Figure 4 compares the instances identified by Panoptic Lifting to 3DIML.
We benchmark all runtimes using a single RTX 3090 post-mask generation. Specifically, comparing their implementation to ours, Panoptic Lifting requires 5.7 hours of training over all scans, with a min and max of 3.6 and 6.6 hours, respectively, since its runtime depends on the number of objects. Contrastive Lifting takes around 3.5 hours on average while 3DIML runs under 20 minutes (14.5 minutes on average) for all scans. Note several components of 3DIML can be easily parallelized, such as dense descriptors extraction using LoFTR and label merging. The runtime of our method is dependent on the number of correspondences produced by LofTR, which doesn’t change for different frontend segmentation models, and we observe similar runtimes for other experiments.
@@
Method Time Panoptic lifting [5] 5.7 hours Contrastive lift [6] 3.6 hours InstanceMap (LoFTR) 6.2 minutes InstanceMap 2.1 minutes InstanceLift (train) 5.6 minutes InstanceLift (refine) 33.2 seconds 3DIML (total) 14.5 minutes
@@
Grounded SAM Table III shows our results for lifting GroundedSAM masks for Replica-vMap. From Figure 5 we see that InstanceLift is effective at interpolating labels missed by InstanceMap and resolving ambiguities produced by GroundedSAM  [^1] GroundedSAM produces lower quality frontend masks than SAM due to prompting using bounding boxes instead of a point grid.. Figure 7 shows that InstanceMap and 3DIML are robust to large viewpoint changes and as well as duplicate objects, assuming nice scans, that is enough context for NetVLAD and LoFTR to somewhat distinguish between them. Table IV and Fig. 6 illustrate our performance on ScanNet [21].
@@
Figure 7: Qualitative results on office3 and room1 from the Replica-vMap split [20]. Both InstanceMap and InstanceLift are able to maintain quality and consistency over the image sequence despite duplicate objects due to sufficient image context overlap across the sequence.